if the product of 2 consecutive numbers is 1332 what are the 2 numbers?
Answer by unlockmath(1688)

Hello,
Let x be one number and (x+1) be the next so now we can write an equation:
x(x+1)=1332 Expand this to:
x^2+x=1332 Subtract 1332 from both sides:
x^2+x-1332=0 Factor this to:
(x-36)(x+37)=0 Solve for x:
x=36
x=-37
Both these will work. So the numbers are 36,37 and -36,-37
Make sense?
RJ
www.math-unlock.com